OpenStudy (anonymous):

minimize f(x,y)=x²+4y² subject to the constraint xy=1 (we are using Langrangian formula) minimize f(x,y)=x²+4y² subject to the constraint xy=1 (we are using Langrangian formula) @Mathematics

OpenStudy (amistre64):

this is quite similar to the other ones right?

OpenStudy (anonymous):

yes but i couldnt solve for all the variables :/

OpenStudy (amistre64):

F(x,y)=x^2+4y^2 ; xy-1=0 F(x,y,L) = x^2+4y^2 + L(xy-1)

OpenStudy (anonymous):

right

OpenStudy (amistre64):

Fx = 2x + Ly 2x + Ly = 0 2x = -Ly x = -Ly/2 Fy = 8y + Lx 8y + Lx = 0 8y + L(-Ly/2) = 0 8y -L^2y/2 = 0 (16y -L^2y)/2 = 0 y(16 -L^2)/2 = 0 y = 2/(16 -L^2) looks messy, but does the math look ok?

OpenStudy (anonymous):

one sec!

OpenStudy (anonymous):

now what?

OpenStudy (amistre64):

im thinking the answer is 4; but that is without going thru the Lagrange mess

OpenStudy (amistre64):

since the constraint is that xy = 1; y = 1/x can be subbed into the original f(x,y)

OpenStudy (amistre64):

other than that, i cant get a good bead on it

OpenStudy (anonymous):

i dont understand :(

OpenStudy (amistre64):

f(x,y)=x²+4y² subject to the constraint xy=1 xy=1 means y = 1/x f(x,y(x))=x^2+4/x^2

OpenStudy (amistre64):

if we do it the other way: f(x,y)=x²+4y² ; xy=1 xy=1 can also mean x = 1/y f(x(y),y)=1/y^2+4y^2

OpenStudy (amistre64):

either way they minimize at 4